# Decay and production properties of strange double charm pentaquark

### v2

@@ III.2 Strong coupling $g_{P_{ccs}\Omega_{cc}\pi}$
Figure 3: The dependence of the strong coupling $g_{P_{ccs}\Omega_{cc}\pi}$ on the Borel mass $M_{B}^{2}$ (left panel) and transfer momentum $Q^{2}$ (right panel ). On the left panel, the transfer momentum is set to be $Q^{2}=m_{\Omega_{cc}}^{2}\sim 13.8\;\mathrm{GeV}^{2}$. On the right panel, the red dots denote the value from Eq. (14) with $s_{0}=22.3\;\mathrm{GeV}^{2}$ and $M_{B}^{2}=1.74\;\mathrm{GeV}^{2}$. The blue solid line is the exponential fitting curve. The two dashed blue lines denote the upper and lower boundary of the uncertainty from various condensates, quark masses and hadronic parameters.
As for $P_{ccs}\Omega_{cc}\pi$ process, in the left panel of Fig. 3, we show the variation of the coupling constant $g_{P_{ccs}\Omega_{cc}\pi}(Q^{2})$ with the Borel mass $M_{B}^{2}$ at $Q^{2}=m_{\Omega_{cc}}^{2}\sim 13.8\;\mathrm{GeV}^{2}$. Such a momentum point is chosen far away from $m_{\pi^{\ast}}^{2}$ so that it can be safely ignored and the OPE series is valid in this region. We find that the coupling constant $g_{P_{ccs}\Omega_{cc}\pi}(Q^{2})$ has a maximum value at $M_{B}^{2}\sim 1.74\;\mathrm{GeV}^{2}$, around which it has minimal dependence on the non-physical parameter $M_{B}^{2}$. We find that the results can be well fitted by the exponential model
$g_{P_{ccs}\Omega_{cc}\pi}(Q^{2})=a\;\mathrm{e}^{-b\;Q^{2}}.$ (22)
The fitting curve is shown in the right panel of Fig. 3, and the result is as follow:
$g_{P_{ccs}\Omega_{cc}\pi}(Q^{2})=-(0.193^{+0.052}_{-0.056}\;\mathrm{GeV}^{-3})\mathrm{e}^{-(0.000\pm 0.00)\;\mathrm{GeV}^{-2}\;Q^{2}}.$ (23)
With the on-shell condition $Q^{2}=-m_{\pi}^{2}$, we can obtain the strong coupling constant $g_{P_{ccs}\Omega_{cc}\pi}(-m_{\pi}^{2})=-0.193$. From the matrix element, we can obtain the decay width for $P_{ccs}^{++}\rightarrow\Xi_{cc}^{++}\bar{K}^{0}$ process:
$\Gamma(P_{ccs}^{++}\rightarrow\Omega_{cc}^{+}\pi^{+})=\frac{\sqrt{\lambda(m_{P_{ccs}}^{2},m_{\Omega_{cc}}^{2},m_{\pi}^{2})}}{8\pi m_{P_{ccs}}^{2}}g_{P_{ccs}\Omega_{cc}\pi}(-m_{\pi}^{2})\left((m_{P_{ccs}}+m_{\Omega_{cc}})^{2}-m_{\pi}^{2}\right).$ (24)
Substitute the above on-shell coupling, we can obtain the decay width as
$\Gamma(P_{ccs}^{++}\rightarrow\Omega_{cc}^{+}\pi^{+})=19.56^{+11.15}_{-10.36}\,\mathrm{MeV}.$ (25)
@@
$\begin{split}\Gamma_{P_{ccs}^{++}}&=\Gamma(P_{ccs}^{++}\rightarrow\Xi_{cc}^{++}\bar{K}^{0})+\Gamma(P_{ccs}^{++}\rightarrow\Omega_{cc}^{+}\pi^{+})\\ &=(84.58^{+19.25}_{-18.80})\,\mathrm{MeV}.\end{split}$ (26)
@@ IV Production via Final-State-Interaction
where the parameters $F_{0}$, $m_{\mathrm{fit}}$ and $\delta$ are taken in Ref. Wang:2017mqp and listed in Tab. 2.
@@
The amplitude for $\Xi_{bc}^{+}\rightarrow D_{s}^{\ast-}\Xi_{cc}^{++}\rightarrow D^{-}P_{ccs}^{++}$ process can be written as:
@@
It should be noted that in some work of final state interaction formalism Chen:2020eyu ; Hsiao:2024szt , the decay amplitude contains the form factor $F(t,m)=(\Lambda^{2}-m_{K}^{2})/(\Lambda^{2}-t)$ for each strong vertices, which is introduced to compensate the off-shell effect of the exchanged particle at the vertices Gortchakov:1995im . In this work, we can contain the off-shell effect with $Q^{2}$-dependent strong coupling, such as $g_{D_{s}^{\ast}DK}(Q^{2})$ and $g_{P_{ccs}\Xi_{cc}\bar{K}}(Q^{2})$. We take the result of strong coupling $g_{D_{s}^{\ast}DK}(Q^{2})$ with QCD sum rule formalism as follow Yang:2024coj :
$g_{D_{s}^{\ast}DK}(Q^{2})=(2.82^{+1.34}_{-0.82}\,\mathrm{GeV^{-2}})\mathrm{e}^{-(0.22\pm 0.00\,\mathrm{GeV}^{-2})Q^{2}}.$ (36)
With strong coupling Eqs. (19), (36) and Eq. (35), and the mass of $\Xi_{bc}^{+}$ taken from the lattice result Brown:2014ena , the decay width of $\Xi_{bc}^{+}\rightarrow P_{ccs}^{++}D^{-}$ process can be calculated as
$\Gamma(\Xi_{bc}^{+}\rightarrow P_{ccs}^{++}D^{-})=(1.17_{-0.40}^{+0.55})\times 10^{-17}\,\mathrm{GeV}.$ (37)
We take the lattice result of lifetime of $\Xi_{bc}^{+}$ Brown:2014ena , the production branching fraction of $\Xi_{bc}^{+}\rightarrow P_{ccs}^{++}D^{-}$ process can be calculate as
$\mathcal{B}r(\Xi_{bc}^{+}\rightarrow P_{ccs}^{++}D^{-})=(4.32_{-1.47}^{+2.02})\times 10^{-6}.$ (38)
@@ V Summary
Based on our previous calculations of mass spectroscopy Yang:2024okq , we further study the decay and production properties of the exotic strange double charm pentaquark state $P_{ccs}^{++}$ with $J^{P}=1/2^{-}$. We perform the three-point QCD sum rules to calculate the coupling constants of $P_{ccs}^{++}\rightarrow\Xi_{cc}^{++}\bar{K}^{0}$ and $P_{ccs}^{++}\rightarrow\Omega_{cc}^{+}\pi^{+}$ strong decay processes. The partial decay widths of these two process are obtained as
$\begin{split}\Gamma(P_{ccs}^{++}\rightarrow\Xi_{cc}^{++}\bar{K}^{0})&=65.02\pm 15.69\,\mathrm{MeV},\\ \Gamma(P_{ccs}^{++}\rightarrow\Omega_{cc}^{+}\pi^{+})&=19.56^{+11.15}_{-10.36}\,\mathrm{MeV},\end{split}$ (39)
@@
$\Gamma_{P_{ccs}^{++}}=84.58^{+19.25}_{-18.80}\,\mathrm{MeV}\,.$ (40)
Furthermore, we study the $\Xi_{bc}^{+}\rightarrow D_{s}^{\ast-}\Xi_{cc}^{++}\rightarrow D^{-}P_{ccs}^{++}$ process via the rescattering mechanism to estimate the branching ratio of $\Xi_{bc}^{+}\rightarrow D^{-}P_{ccs}^{++}$ as $(4.32_{-1.47}^{+2.02})\times 10^{-6}$. This value of branching ratio aligns with the productions of hidden-charm pentaquark states in $\Xi_{b}$ decays Wu:2021caw ; Pan:2023hrk ; Wu:2024lud (with a same $b\rightarrow c\bar{c}s$ weak transitions), which typically around $\mathcal{B}r\sim 10^{-6}-10^{-5}$. Combining with the decay branching ratio $\mathcal{B}r(P_{ccs}^{++}\rightarrow\Xi_{cc}^{++}\bar{K}^{0})$, we obtain $\mathcal{B}r(\Xi_{bc}^{+}\rightarrow D^{-}P_{ccs}^{++}\rightarrow D^{-}\bar{K}^{0}\Xi_{cc}^{++})\approx\mathcal{B}r(\Xi_{bc}^{+}\rightarrow D^{-}P_{ccs}^{++})\mathcal{B}r(P_{ccs}^{++}\rightarrow\Xi_{cc}^{++}\bar{K}^{0})=3.3\times 10^{-6}$, providing a benchmark for future experiments.
To date, LHCb has pursued $\Xi_{bc}^{+}$ state in $\Lambda_{c}\pi$, $\Xi_{c}\pi$ and $DKp$ final states without significant signals LHCb:2020iko ; LHCb:2021xba ; LHCb:2022fbu . Nevertheless, high $\Xi_{bc}^{+}$ yields have been expected at future facilities such as MuIC ($\sim 10^{8}$ events/year) Zhao:2025kvq , CEPC/FCC-ee ($\sim 10^{7}$ events/year) Ma:2025ito , LHeC ($\sim 10^{5}$ events/year) Bi:2017nzv , and LHCb Run 3 ($\sim 10^{4}$ events/year) Zhang:2022jst , producing a considerable amount of double charm pentaquark states in the future.
@@ ACKNOWLEDGMENTS
This work is supported by the National Natural Science Foundation of China with Grant Nos. 12175318, the Natural Science Foundation of Guangdong Province of China under Grant No. 2022A1515011922.
Appendix A Spectrum function for three-point correlation of strong vertices
The spectrum function $\rho(s)$ and $R(M_{B}^{2})$ in Eq. (13) is shown as follow
@@ Appendix A Spectrum function for three-point correlation of strong vertices
The spectrum function $\rho(s)$ and $R(M_{B}^{2})$ in Eq. (14) is shown as follow

### v4

@@ III.2 Strong coupling $g_{P_{ccs}\Omega_{cc}\pi}$
Figure 3: The dependence of the strong coupling $g_{P_{ccs}\Omega_{cc}\pi}$ on the Borel mass $M_{B}^{2}$. The transfer momentum is set to be $Q^{2}=m_{\Omega_{cc}}^{2}\sim 13.8\;\mathrm{GeV}^{2}$.
For the $P_{ccs}\Omega_{cc}\pi$ vertex, we show the variation of the coupling constant $g_{P_{ccs}\Omega_{cc}\pi}(Q^{2})$ with the Borel mass $M_{B}^{2}$ at $Q^{2}=m_{\Omega_{cc}}^{2}\approx 13.8\;\text{GeV}^{2}$ in Fig. 3. Similar to the discussion above, we extract the $1/Q^{2}$ term in the QCD sum rule since the negligible pion mass. The dependence of the coupling constant on the unphysical Borel parameter achieves minimum around $M_{B}^{2}\sim 1.74\;\text{GeV}^{2}$. Neglecting the small $Q^{2}$-dependence, we compute the coupling constant at $Q^{2}=m_{\Omega_{cc}}^{2}$ as
$g_{P_{ccs}\Omega_{cc}\pi}(m_{\Omega_{cc}}^{2})=-(0.19^{+0.05}_{-0.06})\;\text{GeV}^{-3}.$ (21)
From the matrix element, we obtain the decay width for the $P_{ccs}^{++}\to\Omega_{cc}^{+}\pi^{+}$ process:
$\Gamma(P_{ccs}^{++}\rightarrow\Omega_{cc}^{+}\pi^{+})=\frac{\sqrt{\lambda(m_{P_{ccs}}^{2},m_{\Omega_{cc}}^{2},m_{\pi}^{2})}}{8\pi m_{P_{ccs}}^{2}}g_{P_{ccs}\Omega_{cc}\pi}(m_{\Omega_{cc}}^{2})\left((m_{P_{ccs}}+m_{\Omega_{cc}})^{2}-m_{\pi}^{2}\right).$ (22)
Substituting the above coupling, we can obtain the decay width as
$\Gamma(P_{ccs}^{++}\rightarrow\Omega_{cc}^{+}\pi^{+})=20^{+11}_{-10}\,\mathrm{MeV}.$ (23)
@@
$\begin{split}\Gamma_{P_{ccs}^{++}}&=\Gamma(P_{ccs}^{++}\rightarrow\Xi_{cc}^{++}\bar{K}^{0})+\Gamma(P_{ccs}^{++}\rightarrow\Omega_{cc}^{+}\pi^{+})\\ &=(85\pm 19)\,\mathrm{MeV}.\end{split}$ (24)
@@ IV Production via Final-State-Interaction
where the parameters $F_{0}$, $m_{\mathrm{fit}}$ and $\delta$ are taken from Ref. Wang:2017mqp are listed in Tab. 2.
@@
The amplitude for the $\Xi_{bc}^{+}\rightarrow D_{s}^{\ast-}\Xi_{cc}^{++}\rightarrow D^{-}P_{ccs}^{++}$ process can be written as:
@@
It should be noted that in some works of the final state interaction formalism Chen:2020eyu ; Hsiao:2024szt , the decay amplitude contains the form factor $F(t,m)=(\Lambda^{2}-m_{K}^{2})/(\Lambda^{2}-t)$ for each strong vertex, which is introduced to compensate the off-shell effect of the exchanged particle at the vertices Gortchakov:1995im . In this work, we can obtain the strong coupling, such $g_{P_{ccs}\Xi_{cc}\bar{K}}$. We take the result of strong coupling $g_{D_{s}^{\ast}DK}(Q^{2})$ with QCD sum rule formalism as follow Yang:2024coj :
$g_{D_{s}^{\ast}DK}(Q^{2})=(2.8^{+1.3}_{-0.8}\,\mathrm{GeV^{-2}})\mathrm{e}^{-(0.2\pm 0.0\,\mathrm{GeV}^{-2})Q^{2}}.$ (34)
With the values of the strong couplings given in Eqs. (18), (34) and the result of Eq. (33), and the mass of $\Xi_{bc}^{+}$ taken from the lattice result Brown:2014ena , the decay width of $\Xi_{bc}^{+}\rightarrow P_{ccs}^{++}D^{-}$ process can be calculated as
$\Gamma(\Xi_{bc}^{+}\rightarrow P_{ccs}^{++}D^{-})=(1.2_{-0.4}^{+0.6})\times 10^{-17}\,\mathrm{GeV}.$ (35)
Taking the lattice result of lifetime of $\Xi_{bc}^{+}$ Brown:2014ena , the production branching fraction of $\Xi_{bc}^{+}\rightarrow P_{ccs}^{++}D^{-}$ process becomes
$\mathcal{B}r(\Xi_{bc}^{+}\rightarrow P_{ccs}^{++}D^{-})=(4.3_{-1.5}^{+2.0})\times 10^{-6}.$ (36)
@@ V Summary
Based on our previous calculations of mass spectroscopy Yang:2024okq , we further study the decay and production properties of the exotic strange double charm pentaquark state $P_{ccs}^{++}$ with $J^{P}=1/2^{-}$. We perform the three-point QCD sum rules to calculate the coupling constants of $P_{ccs}^{++}\rightarrow\Xi_{cc}^{++}\bar{K}^{0}$ and $P_{ccs}^{++}\rightarrow\Omega_{cc}^{+}\pi^{+}$ strong decay processes. The partial decay widths of these two processes are obtained as
$\begin{split}\Gamma(P_{ccs}^{++}\rightarrow\Xi_{cc}^{++}\bar{K}^{0})&=65\pm 16\,\mathrm{MeV},\\ \Gamma(P_{ccs}^{++}\rightarrow\Omega_{cc}^{+}\pi^{+})&=20^{+11}_{-10}\,\mathrm{MeV},\end{split}$ (37)
@@
$\Gamma_{P_{ccs}^{++}}=85\pm 19\,\mathrm{MeV}\,.$ (38)
Furthermore, we study the $\Xi_{bc}^{+}\rightarrow D_{s}^{\ast-}\Xi_{cc}^{++}\rightarrow D^{-}P_{ccs}^{++}$ process via the rescattering mechanism to estimate the branching ratio of $\Xi_{bc}^{+}\rightarrow D^{-}P_{ccs}^{++}$ as $(4.3_{-1.5}^{+2.0})\times 10^{-6}$. This value of branching ratio aligns with the productions of hidden-charm pentaquark states in $\Xi_{b}$ decays Wu:2021caw ; Pan:2023hrk ; Wu:2024lud (having the same $b\rightarrow c\bar{c}s$ weak transition), which typically lie around $\mathcal{B}r\sim 10^{-6}-10^{-5}$. Combining with the decay branching ratio $\mathcal{B}r(P_{ccs}^{++}\rightarrow\Xi_{cc}^{++}\bar{K}^{0})$, we obtain $\mathcal{B}r(\Xi_{bc}^{+}\rightarrow D^{-}P_{ccs}^{++}\rightarrow D^{-}\bar{K}^{0}\Xi_{cc}^{++})\approx\mathcal{B}r(\Xi_{bc}^{+}\rightarrow D^{-}P_{ccs}^{++})\mathcal{B}r(P_{ccs}^{++}\rightarrow\Xi_{cc}^{++}\bar{K}^{0})=3.3\times 10^{-6}$, providing a benchmark for future experiments.
To date, LHCb has pursued the search of the $\Xi_{bc}^{+}$ state exploring the invariant mass of $\Lambda_{c}\pi$, $\Xi_{c}\pi$ and $DKp$ final states without significant signals LHCb:2020iko ; LHCb:2021xba ; LHCb:2022fbu . Nevertheless, high $\Xi_{bc}^{+}$ yields have been expected at future facilities such as MuIC ($\sim 10^{8}$ events/year) Zhao:2025kvq , CEPC/FCC-ee ($\sim 10^{7}$ events/year) Ma:2025ito , LHeC ($\sim 10^{5}$ events/year) Bi:2017nzv , and LHCb Run 3 ($\sim 10^{4}$ events/year) Zhang:2022jst , which will produce a considerable amount of double charm pentaquark states in the future. Our analysis is specifically applied to the molecular interpretation of the $P_{ccs}^{++}$ state. While compact pentaquark configurations with similar masses are also possible Yang:2024okq , a quantitative analysis of their decay properties within the same QCD sum rule framework proved inconclusive due to the instability of the operator product expansion series for the relevant correlation functions. Future studies with alternative theoretical approaches or improved sum rule constructions may help to clarify the properties of such compact states and their potential distinguishability from molecular candidates in experiments.
@@ ACKNOWLEDGMENTS
This work is supported by the National Natural Science Foundation of China with Grant Nos.  12575153 and 12175318.
Appendix A Spectral function for three-point correlation of strong vertices
The spectral function $\rho(s)$ and $R(M_{B}^{2})$ in Eq. (13) is shown as follow
@@ Appendix A Spectral function for three-point correlation of strong vertices
The spectral function $\rho(s)$ and $R(M_{B}^{2})$ in Eq. (14) is shown as follow